# Clustering in Networks with Time-varying Nodal Attributes

### v1

@@ 3 Learning and Inference
Calculating (7) and the gradient in (8) requires evaluating a conditional expectation under the posterior density $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Here, Langevin dynamics are employed to sample from the posterior distribution, approximating conditional expectations (du2019implicit; nijkamp2020anatomy; purohit2024posterior). In particular, let $k$ be the time step of the Langevin dynamics and let $\delta>0$ be a small step size. The Langevin dynamics to draw samples from the posterior distribution iterate
$\bm{Z}_{i}^{k+1}$ $=\bm{Z}_{i}^{k}+\delta\big[\nabla_{\bm{Z}_{i}}\ln\big(P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i}^{k})\big)-(\bm{Z}_{i}^{k}-\bm{\mu}_{i})\big]+\sqrt{2\delta}\bm{\epsilon}$ (10)
in $k$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ is a random perturbation. The gradient of $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with respect to $\bm{Z}_{i}$ can be calculated through back-propagation. The derivations and computation complexity of the ADMM procedure are provided in our Supplementary Material.
3.1 Nodal Clusterings
After parameter learning, $k$-means method is applied to $\{\hat{\bm{\mu}}_{i}\}$ to cluster the nodes. The LASSO regularization of $\sum_{(i,j)\in\mathcal{E}}\|\bm{\mu}_{i}-\bm{\mu}_{j}\|_{2}$ encourages connected nodes to have similar prior parameters. As a result, the estimated $\{\hat{\bm{\mu}}_{i}\}$ naturally “cluster well” in Euclidean spaces, making the $k$-means method well-suited for partitioning.
The nodal series may contain temporal similarity that are shared by neighboring nodes. By jointly modeling temporal dynamics and structural patterns through the prior parameters, our framework integrates both sources of information to form clusters. Consequently, the resulting partitions are more informative, in the sense that they reflect both temporal similarity and structural proximity, rather than solely relying on the time series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ or the graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})$ alone.
Since most $k$-means algorithms require the number of clusters, $k$, we estimate this parameter via a silhouette score. Specifically, a range of $k$ in $\{2,\dots,K\}$ is considered; the one having the highest silhouette score $S\in[-1,1]$, defined by
@@ 4.1 Simulations
Our simulations generate $50$ graph realizations with a varying number of nodes $N$. The series length $n=100$ is used at every node. To evaluate performance, the following standard clustering metrics are considered: normalized mutual information (NMI), adjusted Rand index (ARI), accuracy score (ACC), homogeneity (HOM), completeness (COM), and cluster purity (PUR). These evaluation metrics require a known truth to measure from; see emmons2016analysis and su2022comprehensive for additional detail. Six competitors, $k$-means, dynamic time warping (DTW) (sakoe2003dtw), covariate assisted spectral clustering (CASC) (binkiewicz2017casc), semi-definite programming (SDP) (yan2021sdp), spectral clustering on network-adjusted covariates (NAC) (hu2024nac), and spectral clustering on ratios-of-eigenvectors (SCORE) (Jin2015SCORE), are provided for comparison.
The $k$-means and DTW methods only use the series for clustering, while the SCORE method only uses graphical structure. The CASC, SDP, and NAC are hybrid methods utilizing both series and graphical structures to cluster nodes, similar to our method. Our simulations assess whether integrating both sources of information leads to better clustering than approaches that use the graph or attributes alone. Performances between our method and existing hybrid approaches are compared. Details of the competitors are provided in the Supplementary Material.
Throughout, our method is dubbed GFL. The latent dimension is taken as $d=3$, and the two hidden layers of the neural network each contain $32$ neurons. The tuning $\lambda$’s are $\{0.1,0.25,0.5,0.75,1.0\}$ and $10\%$ of the series was held out for model selection. For each $\lambda$, $50$ iterations of the ADMM procedure are run and $\gamma=\lambda$ is taken. In each ADMM iteration, the neural network is trained using $20$ steps of an Adam optimizer with a learning rate of $10^{-4}$, while $20$ iterations of block coordinate descent are run for group LASSO. For Langevin dynamic sampling, we take $\delta=0.4$ and $500$ samples are drawn for each node using $50$ MCMC iterations. As the objective function in (2) is highly non-convex due to the neural networks and sub-routines requiring MCMC sampling, specifying standard convergence criteria is impractical. The algorithm is implemented for a preset number of iterations, as an early stopping strategies to avoid over-fitting, which is commonly used in deep learning. The model selection procedure and results for varying latent dimensions as a sensitivity analysis are presented in our Supplementary Material.
@@
Our first scenario simulates graphs with block structures to demonstrate nodal clustering. Specifically, for the node set $\mathcal{V}$ of size $N=|\mathcal{V}|$, the probability matrix $\bm{P}\in[0,1]^{N\times N}$ is defined with entries ${\bm{P}}_{i,j}=0.30\ \forall\ i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\}$, and ${\bm{P}}_{i,j}=0.15$ otherwise. Here, $\mathcal{C}_{1},\mathcal{C}_{2}$, and $\mathcal{C}_{3}$ are three clusters that partition $\mathcal{V}$. Then edges are sampled via an adjacency matrix $\bm{A}\in\{0,1\}^{N\times N}$ format via $\bm{A}_{i,j}\sim\text{Bernoulli}({\bm{P}}_{i,j})$. For $N=120$ nodes, our three unbalanced cluster sizes are $30$, $40$ and $50$; for $N=210$, the cluster sizes are $60$, $70$, and $80$.
@@
$Y_{t,i}=\mu_{k}+\psi_{k}(Y_{t-1,i}-\mu_{k})+\varepsilon_{t,i},\hskip 18.49988ptt=2,\ldots,T,\qquad i\in\mathcal{C}_{k},$
@@
$Y_{1,i}\sim\mathcal{N}\left(\mu_{k},\frac{\sigma_{k}^{2}}{1-\psi_{k}^{2}}\right),\hskip 18.49988pti\in\mathcal{C}_{k}.$ (11)
We choose $\psi_{k}=0.5$ and $\sigma_{k}^{2}=1$ for all three clusters. Cluster means are $\mu_{1}=-1.0$, $\mu_{2}=0.0$, and $\mu_{3}=1.0$. Figure 2 depicts a simulated block graph and its associated series sample means. Table 1 reports sample means and standard deviations of our clustering evaluation metrics over the 50 simulations.
Figure 2: A simulated graph and its time series sample means. The left plot is a simulated graph with unbalanced cluster sizes and $120$ nodes. The right plot shows time series sample means along with the $\pm 1$ standard deviation bands at each time point for the three clusters.
@@
We consider two graph sizes having $N=144$ and $N=196$ nodes, corresponding to $12\times 12$ and $14\times 14$ grids. Both densities are sparser than Scenario 1. The unbalanced cluster sizes are $25$, $30$, $42$, and $47$ for $N=144$ and $35$, $40$, $42$, and $79$ for $N=196$. Figure 3 illustrates a simulated grid and its associated time series means. No clusters of nodes are visually distinguishable from the graph topology (without relying on nodal colors). Table 2 reports the sample means and standard deviations of the evaluation metrics across 50 simulations.
Figure 3: A simulated grid and its associated time series means. The left plot is a simulated graph with 144 nodes having unbalanced cluster sizes. The right plot shows the sample series means of the four clusters; one standard deviation bands are again displayed.
Different from Scenario 1, the grid structure does not reveal cluster boundaries, making clustering based on the graph topology challenging. In this setting, cluster information primarily resides in the nodal time series. The hybrid methods, CASC, SDP, and NAC, which rely on both adjacency matrix and nodal series perform poorly, suggesting that an uninformative spatial structure hinders spectral clustering even when the time series exhibit cluster differences. Moreover, increasing the graph size from $N=144$ to $N=196$ does not improve their performances, and can even degrade matters. Enlarging the grid jeopardizes graph-based approaches that emphasize structural cues. Interestingly, the SCORE method, which does not utilize any series for clustering, performs well when the correct number of clusters is supplied. The slight error here stems from the method’s tendency to produce four equal-sized clusters, making some boundary nodes incorrectly assigned. In contrast, our GFL method continues to perform nicely, adapting to the absence of cluster signals in graph and relying on informative nodal series.
@@
Scenario 3 extends Scenario 1 by including correlation within series in the same cluster. Here, a VAR model of order one is applied to incorporate cluster-specific correlations. The $N$-dimensional time series $\{{{\bm{Y}}_{t}}\}_{t=1}^{T}$ are generated simultaneously over all nodes via
${\bm{Y}}_{t}={\bm{\beta}}+{\Phi}({\bm{Y}}_{t-1}-{\bm{\beta}})+{\bm{\xi}}_{t},\qquad t=2,\ldots,T$
where $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{N})^{\top}$ is the nodal mean vector: $\beta_{i}=\mu_{k}$ for $i\in\mathcal{C}_{k}$. Here, the error $\bm{\xi}_{t}$ is sampled as $\mathcal{N}({\bm{0}}_{N},\bm{\Sigma})$ for each $t$, where $\bm{\Sigma}=\{\Sigma_{i,j}\}_{i,j=1,\ldots,N}$ has the intra-cluster correlations $\Sigma_{i,j}=1\ \forall\ i=j$, $\Sigma_{i,j}=\rho\ \forall\ i\neq j,\text{ and }i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\}$, and $\Sigma_{i,j}=0$ otherwise. We take $\rho=0.3$, which induces moderate correlation.
The VAR coefficient is taken as $\Phi=0.5\bm{I}_{N}$ and the means for the three clusters are repeated from Scenario 1: $\mu_{1}=-1.0$, $\mu_{2}=0.0$, and $\mu_{3}=1.0$. Series from different clusters are statistically independent. Our simulations start with ${\bm{Y}}_{1}={\bm{0}}$ and burn-in is used for $100$ iterations to ensure stationary nodal series. The same three cluster block graph in Scenario 1 is used. Figure 4 shows a simulated grid graph and the associated series sample means. Comparing to Scenario 1, the series means are not as “well separated”, making clustering more difficult.
Table 3 reports the means and standard deviations of our evaluation metrics over 50 simulations. With $120$ nodes, $k$-means and CASC methods again provide good clusterings. CASC clusterings, which leverage nodal covariates via pair-wise similarity, benefits from the positive intra-cluster correlations. When the number of nodes increases to $210$, competitors improve significantly (more than in Scenario 1), exploiting the temporal correlation between nodes to improve performance. Our GFL method slightly degrades from Scenario 1, but still maintains accuracy and is often the best method. Even in the presence of modest intra-nodal correlation, our GFL clustering method continues to perform well.
@@ 4.2 California County Temperatures
We now cluster counties in the state of California via observed temperatures. There are $N=58$ counties in the state, all of which report data. Monthly average temperatures over the 14-year period January 2011 - December 2024 are studied. The data come from the National Centers for Environmental Information (download details are listed below). Recent clusterings (by station, not county) of California climate appear in abatzoglou2009classification. While we examine the 14 year record to reduce impact of trends on results, California climate trend studies appear in ladochy2007recent and walton2018assessment.
A network with 58 nodes representing all counties was generated; two counties are connected by an edge if they share some common border. The goal here is to utilize both spatial proximity of the counties and their temperature series to identify useful climate regions within the state. Some pre-processing was implemented to obtain stationary series. Specifically, with $\{Y_{t}\}_{t=1}^{n}$ denoting a county monthly series, we convert to a standardized measurement by examining
@@
$\hat{\mu}_{\nu}=\frac{1}{n_{\rm yr}}\sum_{k=0}^{n_{\rm yr}-1}Y_{kT+\nu},\qquad\qquad\hat{\sigma}_{\nu}^{2}=\frac{\sum_{k=0}^{n_{\rm yr}-1}(Y_{kT+\nu}-\hat{\mu}_{\nu})^{2}}{n_{\rm yr}-1}.$
This standardization removes the monthly mean and variance seasonality of temperatures, but preserves county-level variability. In this analysis, the latent dimension $d$ is taken as three. Sensitivity analyses from the simulation study, provided in the Supplementary Material, suggest that this choice is adequate. Figure 5(a) displays the results from our methods. Ten clusters were selected by silhouette scores.
(a) GFL
(b) $k$-means ($k=10$)
Figure 5: California county clustering from GFL and $k$-means.
The clustering is sensible, placing desert Southeast, Mid-Coastal, Lower Coastal, Bay Area, Sierra Nevada Mountain, San Joaquin Valley, Coastal Northwest, and Mountainous Northeast counties in distinct clusters. California has many micro-climates and clustering its counties is challenging.
For comparison, Figure 5(b) shows a $k$-means clustering, the best competitor in simulation study, with ten clusters. Ten clusters were chosen to be comparable to our above clustering (if the number of clusters is selected by silhouette scores, fewer clusters and unrealistic results arise). This clustering is not as pleasing. Specifically, Inyo County in the Owens Valley has been placed in the desert south cluster (as have the coastal counties northwest of Los Angeles). The Northeast Mountain counties now mix with the Coastal Northwest counties. With understanding of the local geography, these are implausible.

### v2

@@ 3 Learning and Inference
Calculating (6) and the gradient in (7) requires evaluating a conditional expectation under the posterior density $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Here, Langevin dynamics are employed to sample from the posterior distribution, approximating conditional expectations (Du and Mordatch, 2019; Nijkamp et al., 2020; Purohit et al., 2024). In particular, let $k$ be the time step of the Langevin dynamics and let $\delta>0$ be a small step size. The Langevin dynamics to draw samples from the posterior distribution iterate
$\begin{split}\bm{Z}_{i}^{k+1}=\bm{Z}_{i}^{k}+\delta\big[&\nabla_{\bm{Z}_{i}}\ln\big(P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i}^{k})\big)\\ &-(\bm{Z}_{i}^{k}-\bm{\mu}_{i})\big]+\sqrt{2\delta}\bm{\epsilon}\end{split}$ (9)
in $k$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ is a random perturbation to the process. The gradient of $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with respect to the latent variable $\bm{Z}_{i}$ is calculated through back-propagation. The derivations of the updates and computation complexity of the ADMM procedure are provided in the Appendix.
4 Nodal Clustering and Model Selection
4.1 Nodal Clustering
For graphs with dynamic attributes, time series observed at neighboring nodes often exhibit similar evolving patterns. To capture this behavior, the proposed framework jointly models temporal dynamics and graphical structures, allowing the learned priors to integrate both sources of information. Specifically, the graph-fused LASSO regularization encourages adjacent nodes with similar temporal behavior to have similar prior means, effectively promoting piecewise-constant structure over the graph. The estimated prior means form well-separated regions in a low-dimensional Euclidean space. Therefore, $k$-means method is applied to $\{\hat{\bm{\mu}}_{i}\}_{i\in\mathcal{V}}$, as centroid-based clustering method align well with the geometry induced by the regularization. The resulting partitions are more informative, in the sense that they reflect both temporal similarity and structural proximity, rather than solely relying on the time series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ or the graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})$ alone.
Since $k$-means algorithm requires the number of clusters $k$, we estimate this parameter via a silhouette score. A range of $k$ in $\{2,\dots,K\}$ is considered; the one having the highest silhouette score $S\in[-1,1]$, defined by
@@ 4.1 Nodal Clustering
4.2 Model Selection
The optimization problem in (3) involves the tuning parameter $\lambda$. Given a list of candidate $\lambda$ values, model selection is performed through cross validation. Specifically, the observed data is split into training and testing sets via nodes. Let $\mathcal{V}_{\text{test}}$ be the node set of the testing data. During training, we replace the node $i$ data $\bm{Y}_{i}$ with the element-wise average $\bar{\bm{Y}}_{i}$ from all neighbors. The node $i$ data is held out for testing when $i\in\mathcal{V}_{\text{test}}$.
$\sum_{i\in\mathcal{V}_{\text{test}}}\ln\big(P(\bm{Y}_{i})\big)\approx\sum_{i\in\mathcal{V}_{\text{test}}}\ln\left[\frac{1}{s}\sum_{u=1}^{s}\big[P(\bm{Y}_{i}|\bm{Z}_{u,i})\big]\right]$
@@ 5.1 Simulations
Our simulations generate $50$ graph realizations with a varying number of nodes $N\coloneqq|\mathcal{V}|$. The series length $n=100$ is used at every node. To evaluate performance, the following standard clustering metrics are considered: normalized mutual information (NMI), adjusted Rand index (ARI), accuracy score (ACC), homogeneity (HOM), completeness (COM), and cluster purity (PUR). These evaluation metrics require a known truth to measure from; see (Emmons et al., 2016) and (Su et al., 2022) for additional details. Six competitors, $k$-means, dynamic time warping (DTW) (Sakoe and Chiba, 2003), covariate assisted spectral clustering (CASC) (Binkiewicz et al., 2017), semi-definite programming (SDP) (Yan and Sarkar, 2021), spectral clustering on network-adjusted covariates (NAC) (Hu and Wang, 2024), and spectral clustering on ratios-of-eigenvectors (SCORE) (Jin, 2015), are provided for comparison.
The $k$-means and DTW methods only use the series for clustering, while the SCORE method only uses graphical structure. The CASC, SDP, and NAC are hybrid methods utilizing both series and graphical structures to cluster nodes, similar to our method. Our simulations assess whether integrating both sources of information leads to better clustering than approaches that use the graph or attributes alone; then performances against the hybrid approaches are compared. Details of the competitors are provided in the Appendix.
Throughout, our method is dubbed GFL. The latent dimension is taken as $d=3$, and the two hidden layers of the neural network each contain $32$ neurons. The tuning $\lambda$’s are $\{0.1,0.25,0.5,0.75,1.0\}$ and $10\%$ of the series was held out for model selection. For each $\lambda$, $50$ iterations of the ADMM procedure are run and $\gamma=\lambda$ is taken. In each ADMM iteration, the neural network is trained using $20$ steps of an Adam optimizer with a learning rate of $10^{-4}$. For Langevin dynamic sampling, we take $\delta=0.4$ and $500$ samples are drawn for each node using $50$ MCMC iterations. The results for varying latent dimensions as a sensitivity analysis are presented in the Appendix.
@@
Our first scenario simulates graphs with block structures to demonstrate nodal clustering. Specifically, for the node set $\mathcal{V}$ of size $N\coloneqq|\mathcal{V}|$, a probability matrix $\bm{P}\in[0,1]^{N\times N}$ is first defined as
$\bm{P}_{ij}=\begin{cases}0.30,&\ i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\},\\ 0.15,&\ \text{otherwise,}\end{cases}$
where $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ are three clusters that partition the node set $\mathcal{V}$. Then edges are sampled via an adjacency matrix format as $\bm{A}_{i,j}\sim\text{Bernoulli}({\bm{P}}_{i,j})$. For $N=120$ nodes, our three unbalanced cluster sizes are $30$, $40$ and $50$; for $N=210$, the cluster sizes are $60$, $70$, and $80$.
@@
$Y_{t,i}=\mu_{k}+\psi_{k}(Y_{t-1,i}-\mu_{k})+\varepsilon_{t,i},\quad t=2,\ldots,T,\quad i\in\mathcal{C}_{k},$
@@
$Y_{1,i}\sim\mathcal{N}\left(\mu_{k},\frac{\sigma_{k}^{2}}{1-\psi_{k}^{2}}\right),\qquad i\in\mathcal{C}_{k}.$ (10)
We choose $\psi_{k}=0.5$ and $\sigma_{k}^{2}=1$ for all three clusters. Cluster means are $\mu_{1}=-1.0$, $\mu_{2}=0.0$, and $\mu_{3}=1.0$. Figure 2 depicts a simulated block graph and its associated series means. Table 1 reports sample means and standard deviations of our clustering evaluation metrics.
Figure 2: A simulated graph and its time series sample means. The left plot is a simulated graph with unbalanced cluster sizes and $120$ nodes. The right plot shows time series sample means along with the $\pm 1$ standard deviation bands at each time point for the clusters.
@@
We consider two graph sizes having $N=144$ and $N=196$ nodes. Both edge densities are sparser than Scenario 1. The unbalanced cluster sizes are $25$, $30$, $42$, and $47$ for $N=144$ and $35$, $40$, $42$, and $79$ for $N=196$. Figure 3 illustrates a simulated grid and its associated time series means. No clusters of nodes are visually distinguishable from the graph topology (without relying on nodal colors). Table 2 reports the sample means and standard deviations of the evaluation metrics across $50$ simulations.
Figure 3: A simulated grid and its associated time series means. The left plot is a simulated graph with 144 nodes having unbalanced cluster sizes. The right plot shows the sample series means of the four clusters; one standard deviation bands are displayed.
Different from Scenario 1, the grid structure does not reveal cluster boundaries, making clustering based on the graph topology challenging. In this setting, cluster information resides in the nodal time series. The hybrid methods, CASC, SDP, and NAC, which rely on both adjacency matrix and nodal series perform poorly, suggesting that an uninformative spatial structure hinders spectral clustering even when the time series exhibit cluster differences. Moreover, increasing the graph size $N$ from $144$ to $196$ does not improve their performances, and can even degrade matters. Enlarging the grid jeopardizes graph-based approaches that emphasize structural cues. Interestingly, the SCORE method, which does not utilize any series for clustering, performs well when the correct number of clusters is supplied. The slight error here stems from the method’s tendency to produce four equal-sized squared regions, making some boundary nodes incorrectly assigned. In contrast, our GFL method continues to perform nicely, adapting to the absence of cluster signals in graph and relying on informative nodal series.
@@
Scenario 3 extends the block graphs in Scenario 1 by including correlations within the series in the same cluster. Here, a VAR model of order one is applied to incorporate cluster-specific correlations. The $N$-dimensional time series $\{{{\bm{Y}}_{t}}\}_{t=1}^{T}$ are generated simultaneously over all nodes via
${\bm{Y}}_{t}={\bm{\beta}}+{\Phi}({\bm{Y}}_{t-1}-{\bm{\beta}})+{\bm{\xi}}_{t},\quad t=2,\ldots,T$
where $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{N})^{\top}$ is the nodal mean vector with $\beta_{i}=\mu_{k}$ for $i\in\mathcal{C}_{k}$. The error $\bm{\xi}_{t}$ is sampled as $\mathcal{N}({\bm{0}}_{N},\bm{\Sigma})$ for each $t$, where $\bm{\Sigma}=\{\Sigma_{i,j}\}_{i,j=1,\ldots,N}$ has the intra-cluster correlations $\Sigma_{i,j}=1\ \forall\ i=j$, $\Sigma_{i,j}=\rho\ \forall\ i\neq j,\text{ and }i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\}$, and $\Sigma_{i,j}=0$ otherwise. We take $\rho=0.3$, which induces moderate correlation.
The VAR coefficient is taken as $\Phi=0.5\bm{I}_{N}$ and the means for the three clusters are repeated from Scenario 1. Series from different clusters are statistically independent. Our simulations start with ${\bm{Y}}_{1}={\bm{0}}$ and burn-in is used for $100$ iterations to ensure stationary nodal series. Figure 4 shows a simulated block graph and the associated series sample means. Comparing to Scenario 1, the series means are not as well separated, making clustering more difficult.
Table 3 reports the means and standard deviations of our evaluation metrics over 50 simulations. With $120$ nodes, $k$-means and CASC methods again provide good clusterings. CASC, which leverages nodal covariates via pair-wise similarity, benefits from the positive intra-cluster correlations. When the number of nodes increases to $210$, competitors improve significantly (more than in Scenario 1), exploiting the temporal correlation between nodes to improve performance. Our GFL method slightly degrades from Scenario 1, but still maintains accuracy and is often the best method. Even in the presence of modest intra-nodal correlation, our GFL clustering method continues to perform well.
@@ 5.2 California County Temperatures
We now cluster counties in the state of California via observed temperatures. California contains both the highest and lowest elevations in the contiguous United States. Its coastal proximity to the Pacific Ocean, neighboring desert environments, and broad latitudinal span lead to significant temperature heterogeneity, making clustering counties in California an interesting yet challenging problem. There are $N=58$ counties in the state, all of which report data. Monthly average temperatures over the $14$-year period January 2011 - December 2024 are studied. The data come from the National Centers for Environmental Information. Recent clusterings (by station, not county) of California climate appear in (Abatzoglou et al., 2009). While we examine the $14$-year record to mitigate the impact of non-stationarity on results, California climate trend studies appear in (LaDochy et al., 2007) and (Walton and Hall, 2018).
A network with $58$ nodes representing all counties was generated; two counties are connected by an edge if they share common border. The goal here is to utilize both spatial proximity of the counties and their temperature series to identify useful climate regions within the state. Some pre-processing was implemented to obtain stationary series. Specifically, with $\{Y_{t}\}_{t=1}^{n}$ denoting a county monthly series, we convert to a standardized measurement by examining
@@
$\hat{\mu}_{\nu}=\frac{1}{n_{\rm yr}}\sum_{k=0}^{n_{\rm yr}-1}Y_{kT+\nu},\quad\hat{\sigma}_{\nu}^{2}=\frac{\sum_{k=0}^{n_{\rm yr}-1}(Y_{kT+\nu}-\hat{\mu}_{\nu})^{2}}{n_{\rm yr}-1}.$
This standardization removes the monthly mean and variance seasonality of temperatures, but preserves county-level variability. In this analysis, the latent dimension $d$ is taken as three. Sensitivity analyses from the simulation study, provided in the Appendix, suggest that this choice is adequate. Figure 5 displays the results from our methods. Ten clusters were selected by silhouette scores. Results from competitors are discussed in the Appendix.
Figure 5: California county clustering from GFL.
The separation between coastal and inland regions is apparent, and the latitudinal transition from north to south is also observed. Cluster 2 and Cluster 8 are dominated by mountainous areas, whereas Clusters 1, 3, 7, 9 are characterized by lower elevations. The Bay Area is further divided into a Delta region (Cluster 1) and coastal region (Cluster 10), indicating that the clustering method has captured intricate climatic heterogeneity within the area. Moreover, Fresno, Tulare, and Kern Counties are leading contributors to agricultural production in the United States, and they are grouped in Cluster 5 at the southern area of the Central Valley, a basin that receives upstream river inflow with sufficient heat. Despite counties are administratively defined, the resulting clusters exhibit clear regional differentiation.
In summary, the clustering is sensible, placing Desert Southeast, Mid-Coastal, Lower Coastal, Bay Area, San Joaquin Valley, Sierra Nevada Mountain, Coastal Northwest, and Mountainous Northeast counties into distinct regions. While the clustering is based on temperature dynamics, the resulting regions reveal underlying geographic and agricultural characteristics, with potential implications for resource allocation and policy planning.